# IFT-UAM/CSIC-25-108

### v1

@@ 3.1 Gauge ensembles and chiral trajectory
$\beta$
$a$
$\mathrm{fm}]$
id $L/a$ $T/a$
$m_{\pi}$
$\mathrm{MeV}]$
$m_{K}$
$\mathrm{MeV}]$
$m_{\pi}L$ LMD MDU $N_{\mathrm{cnfg}}$
@@ 3.4 Matching procedure of sea and valence sectors
These expressions correspond to polynomial approximations in the valence mass parameters ($\kappa^{\mathrm{\scriptscriptstyle(val)}}$, $a\mu_{1}$, $a\mu_{3}$) around the matching point, denoted by the target parameters $\left(\kappa^{\mathrm{\scriptscriptstyle(val)}*},a\mu_{1}^{*},a\mu_{3}^{*}\right)$. A combined fit of $\phi_{2}^{\mathrm{Wtm}}$, $\phi_{4}^{\mathrm{Wtm}}$, and $m_{12}^{\mathrm{Wtm}}$ — as given in Eqs. (3.78)–(3.80) — is performed for each ensemble. This approach exploits the common dependence on the parameters $\kappa^{\mathrm{val}*}$, $a\mu_{1}^{*}$, and $a\mu_{3}^{*}$, in addition to the individual dependence on the parameters $p_{i}$ for $i=1,\dots,8$. The matching and maximal twist tuning procedure is illustrated in Fig. 6.  [^9] In addition to the r.h.s. of Eqs. (3.78)–(3.80), further functional forms have been explored. However, it was observed that the target matching parameters remain insensitive to these variations, given that only minor interpolations are involved in this particular step of the analysis.
@@ 4.1 Determination of $t_{0}^{\mathrm{ph}}$ from $f_{\pi K}^{\mathrm{isoQCD}}$
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1444(10)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.117)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1442(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (4.118)
@@ 4.2 Determination of $t_{0}^{\mathrm{sym}}$ and the lattice spacing $a$
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1441(10)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.126)
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1439(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,$, (4.127)
@@ 4.3 Determination of $t_{0}^{*}$
$\sqrt{t_{0}^{*}}$ $=0.1437(9)(4)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (4.132)
$\sqrt{t_{0}^{*}}$ $=0.1439(9)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.133)
$\sqrt{t_{0}^{*}}$ $=0.1439(7)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (4.134)
@@ 4.4 Determination of $t_{0}^{\mathrm{ph}}$ from $f_{\pi}^{\mathrm{isoQCD}}$
$\Phi_{K,\chi\mathrm{SU(3)}}^{\mathrm{cont}}(\phi_{2})\equiv\left(\sqrt{8t_{0}}f_{K}\right)^{\mathrm{cont}}$ $=\frac{A}{4\pi}\left[1+\frac{B}{2}\left(\phi_{4}-\phi_{2}\right)+C\phi_{4}\right$.
@@ 5 Conclusions
$\sqrt{t_{0}^{\mathrm{ph}}}=0.1442(6)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.145)
@@
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1439(6)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.146)
$\sqrt{t_{0}^{*}}$ $=0.1439(7)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.147)
@@ Acknowledgements
We acknowledge PRACE for awarding us access to MareNostrum at Barcelona Supercomputing Center (BSC), Spain and to HAWK at GCS@HLRS, Germany. The authors thankfully acknowledge the computer resources at MareNostrum and the technical support provided by Barcelona Supercomputing Center (FI-2020-3-0026). We thank CESGA for granting access to the FinisTerrae systems. This work is partially supported by grants PGC2018-094857-B-I00, PID2021-127526NB-I00, and PID2024-160152NB-I00, funded by MCIN/AEI/10.13039/501100011033, by “ERDF A way of making Europe”, and by FEDER, UE, as well as by the Spanish Research Agency (Agencia Estatal de Investigación) through grants IFT Centro de Excelencia Severo Ochoa SEV-2016-0597 and No. CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033, and by the European Commission – NextGenerationEU, through Momentum CSIC Programme: Develop Your Digital Talent. We also acknowledge support from the project H2020-MSCAITN-2018-813942 (EuroPLEx), under grant agreement No. 813942, and the EU Horizon 2020 research and innovation programme, STRONG-2020 project, under grant agreement No. 824093. This work is partially funded by the European Commission - NextGenerationEU, through Momentum CSIC Programme: Develop Your Digital Talent. We acknowledge HPC support by Emilio Ambite, staff hired under the Generation D initiative, promoted by Red.es, an organisation attached to the Spanish Ministry for Digital Transformation and the Civil Service, for the attraction and retention of talent through grants and training contracts, financed by the Recovery, Transformation and Resilience Plan through the European Union’s Next Generation funds.
@@ B.5 Automatic $\mbox{O}(a)$-improvement and the partially quenched mixed-action setup
We start by considering the argument for automatic $\mbox{O}(a)$-improvement at maximal twist, first put forward in Ref. [13].  [^12] See also Refs. [98, 99, 100, 38, 101] for streamlined versions of the argument and further insights into the structure of automatic $\mbox{O}(a)$-improvement. In the fully twisted case, subtracted standard quark masses are set to zero, possibly up to $\mbox{O}(a)$ effects. Consequently, all terms proportional to $\mathbf{m}$ in Eqs. (B.214)–(B.215) vanish. Equivalently, the terms depending on (combinations of) $m_{i}$ vanish in Eqs. (B.216)–(B.220) and (B.223)–(B.227). Concerning the terms proportional to (combinations of) $\mu_{i}$, it is straightforward to verify that they systematically possess opposite physical parity to the dimension-three fields they correct. While parity is not preserved when $\boldsymbol{\mu}\neq 0$, it is broken only at $\mbox{O}(a)$, which means that the matrix elements of parity-odd operators will be $\mbox{O}(a)$. As a consequence, the insertion of any of these counterterms in on-shell correlation functions will give rise to terms of $\mbox{O}(a^{2})$, since the $\mbox{O}(a)$ amplitude will be multiplied by the explicit factor of $a$ that lowers the field dimension. It follows that no Symanzik counterterms to the currents are needed in order to obtain physical observables that scale towards the continuum as $\mbox{O}(a^{2})$. An equivalent conclusion, as discussed in Ref. [95], can also be obtained using the Wilson Averaging (WA) procedure introduced in Ref. [13], which involves averaging correlation functions with opposite signs of the Wilson parameter $r_{i}$. The discrete symmetries of the Wtm regularisation described in Sect. B.2 allow the WA to be reformulated as a Mass Average (MA), where correlation functions with opposite signs of the subtracted mass $m_{i}$ are considered. The automatic $\mbox{O}(a)$-improvement of physical observables at maximal twist follows from the limiting case in which the MA is performed with $m_{i}=0$ [95]. The argument remains valid when the Sheikholeslami-Wohlert term [9] is included, as its contribution to physical observables in the fully twisted case is of order $\mbox{O}(a^{2})$. Moreover, employing a non-perturbative determination of the improvement coefficient $c_{\mathrm{sw}}$ [102, 35] eliminates double insertions of the Pauli term, also contributing at $\mbox{O}(a^{2})$, within the SymEFT framework.
@@
We now turn to the considered mixed-action formulation, corresponding to a partially quenched setup in which the twisted-mass parameters appear exclusively in the valence sector. A first observation can be made when considering quenching. If contributions from sea quarks to the dynamics are neglected, then the terms in Eqs. (B.214)–(B.215) originating from sea-quark loops vanish. These terms can be readily identified as those containing factors of ${\mathrm{tr}}(\mathbf{m})$ and ${\mathrm{tr}}(\boldsymbol{\mu})$, i.e., those associated with improvement coefficients $\bar{b}_{\scriptscriptstyle\Gamma}$ and $\hat{b}_{\scriptscriptstyle\Gamma}$, respectively. For example, in the case of quenched standard Wilson fermions, only counterterms with coefficients $c_{\Gamma}$, $b_{\scriptscriptstyle\Gamma}$, and $f_{\scriptscriptstyle\Gamma}$ need to be considered. Conversely, when present, the terms generated solely by quark loops have coefficients that, in a perturbative expansion, start at order $\mbox{O}(\alpha_{\rm\scriptscriptstyle s}^{2}(1/a))$. This applies to both coefficients $\bar{b}_{\scriptscriptstyle\Gamma}$ and $\hat{b}_{\scriptscriptstyle\Gamma}$.
@@
As noted above, the $\mbox{O}(a)$ terms that are not eliminated depend on the improvement coefficients $\bar{b}_{\Gamma}$ which arise only at order $\mbox{O}(\alpha_{\rm\scriptscriptstyle s}^{2})$. Moreover, for the considered $N_{\mathrm{\scriptstyle f}}=2+1$ CLS ensembles, the light and strange sea-quark masses satisfy $a\mathrm{tr}\left(\mathbf{m}^{\mathrm{\scriptscriptstyle(sea)}}\right)\lesssim 10^{-2}$. Overall, the $a\bar{b}_{\Gamma}\mathrm{tr}\left(\mathbf{m}^{\mathrm{\scriptscriptstyle(sea)}}\right)$ term is therefore expected to introduce only a very small correction. In the present study, this type of residual contributions are neglected in both the unitary and mixed-action setups. The dedicated numerical studies reported in Sec. 4 indicate that the continuum-limit scaling of the physical quantities based on these two regularisations are consistent with leading lattice artefacts of $\mbox{O}(a^{2})$, as can be seen, for instance, in Fig. 8.
@@ E.1 Statistical error analysis
The choice of $W$ necessitates careful consideration [110, 111]. Values that are too small risk introducing systematic uncertainties through truncation of the tail of the autocorrelation function, whereas excessively large values can lead to inflated statistical fluctuations by incorporating data from regions in which the signal is lost. The exponential decay of $\overline{\Gamma}_{F}^{\alpha}(t)$ at asymptotically large $t$ is governed by the exponential autocorrelation time $\tau_{\mathrm{exp}}^{\alpha}$, corresponding to the slowest mode of the Markov process. The corresponding truncation error is therefore of order $\mbox{O}(\exp(-W/\tau_{\mathrm{exp}}^{\alpha}))$. Following the proposal of Ref. [110], we select the value $W=\overline{W}$ that minimises the combined statistical and systematic error, under the assumption $\tau_{\mathrm{exp}}^{\alpha}\approx S_{\tau}\tau_{\mathrm{int}}^{\alpha}$. The default value of the parameter $S_{\tau}$ is set to $S_{\tau}=4$. Direct inspection of the stability of $\tau_{\mathrm{int}}^{\alpha}(F)$ as a function of $W$ is used to adjust the value of $S_{\tau}$ until $\overline{W}$ is observed to lie within a plateau regime.
@@ I Variation in chi-squared weights and model-selection criteria
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1444(10)(5)
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1442(6)(4)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1446(9)(5)
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1444(5)(4)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1444(10)(5)[11]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1442(6)(4)[7]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1444(8)(5)[9]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1442(6)(3)[7]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1446(9)(5)[10]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1443(5)(4)[6]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1442[10]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1441(8)(4)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (J.288)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1438(10)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (J.289)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1441(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (J.290)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1444(9)(6)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (J.292)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1451(9)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (J.293)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1449(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (J.294)
@@
Figure 33: Dependence of the value of $\sqrt{t_{0}^{\mathrm{ph}}}$ on the prescription adopted to define the isoQCD limit. The prescription adopted in this work follows the Edinburgh consensus, as defined in the FLAG 24 report [2], and presented in Eqs. (4.83)–(4.85), supplemented by the isoQCD value of $f_{K}^{\mathrm{isoQCD}}$ [2], as given in Eq. (4.86). The alternative schemes in Eqs. (J) and (J) are based on the FLAG 21 [1] and FLAG 16 [124] reports, respectively. The $N_{\mathrm{\scriptstyle f}}=2+1$ FLAG 24 average result [2] is denoted by the black symbol and the vertical band.
The results obtained with different isoQCD prescriptions can be mapped onto one another by using the derivatives of the relevant observables with respect to the scheme-defining quantities. Table 10 lists the partial derivatives of $\sqrt{t_{0}^{\mathrm{ph}}}$ with respect to $m_{\pi,K}$ and $f_{\pi,K}$, evaluated at their isoQCD values specified in Eqs. (4.83)–(4.86). We have explicitly verified that the shifts obtained from the derivatives in Table 10 reproduce the values quoted in Eqs. (J.288) and (J.294) – i.e. the results for the FLAG 21 and FLAG 16 isoQCD prescriptions, respectively – when starting from Eq. (4.118) which is based on the FLAG 24 input values in Eqs. (4.83)–(4.86).

### v2

@@ 3.1 Gauge ensembles and chiral trajectory
$\beta$ $a$ $\mathrm{fm}]$ id $L/a$ $T/a$ $m_{\pi}$ $\mathrm{MeV}]$ $m_{K}$ $\mathrm{MeV}]$ $m_{\pi}L$ LMD MDU $N_{\mathrm{cnfg}}$
@@ 3.4 Matching procedure of sea and valence sectors
[…]
@@ 4.1 Determination of $t_{0}^{\mathrm{ph}}$ from $f_{\pi K}^{\mathrm{isoQCD}}$
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1441(8)(4)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.117)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1440(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (4.118)
@@ 4.2 Determination of $t_{0}^{\mathrm{sym}}$ and the lattice spacing $a$
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1438(10)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.126)
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1437(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,$, (4.127)
@@ 4.3 Determination of $t_{0}^{*}$
$\sqrt{t_{0}^{*}}$ $=0.1435(9)(4)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (4.132)
$\sqrt{t_{0}^{*}}$ $=0.1438(9)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (4.133)
$\sqrt{t_{0}^{*}}$ $=0.1437(7)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (4.134)
@@ 4.4 Determination of $t_{0}^{\mathrm{ph}}$ from $f_{\pi}^{\mathrm{isoQCD}}$
$\Phi_{K,\chi\mathrm{SU(3)}}^{\mathrm{cont}}(\phi_{2})\equiv\left(\sqrt{8t_{0}}f_{K}\right)^{\mathrm{cont}}$ $=\frac{A}{4\pi}\left[1+B\left(\phi_{4}-\frac{1}{2}\phi_{2}\right)+C\phi_{4}\right$.
@@ 5 Conclusions
$\sqrt{t_{0}^{\mathrm{ph}}}=0.1440(6)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.145)
@@
$\sqrt{t_{0}^{\mathrm{sym}}}$ $=0.1437(6)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.146)
$\sqrt{t_{0}^{*}}$ $=0.1437(7)(4)\;\mathrm{fm}\;\left[f_{\pi K}\right]\,$, (5.147)
@@ Acknowledgements
We acknowledge PRACE for awarding us access to MareNostrum at Barcelona Supercomputing Center (BSC), Spain and to HAWK at GCS@HLRS, Germany. The authors thankfully acknowledge the computer resources at MareNostrum and the technical support provided by Barcelona Supercomputing Center (FI-2020-3-0026). We thank CESGA for granting access to the FinisTerrae systems. This work is partially supported by grants PGC2018-094857-B-I00, PID2021-127526NB-I00, and PID2024-160152NB-I00, funded by MCIN/AEI/10.13039/501100011033, by “ERDF A way of making Europe”, and by FEDER, UE, as well as by the Spanish Research Agency (Agencia Estatal de Investigación) through grants IFT Centro de Excelencia Severo Ochoa SEV-2016-0597 and No. CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033, and by the European Commission – NextGenerationEU, through Momentum CSIC Programme: Develop Your Digital Talent. We also acknowledge support from the project H2020-MSCAITN-2018-813942 (EuroPLEx), under grant agreement No. 813942, and the EU Horizon 2020 research and innovation programme, STRONG-2020 project, under grant agreement No. 824093. We acknowledge HPC support by Emilio Ambite, staff hired under the Generation D initiative, promoted by Red.es, an organisation attached to the Spanish Ministry for Digital Transformation and the Civil Service, for the attraction and retention of talent through grants and training contracts, financed by the Recovery, Transformation and Resilience Plan through the European Union’s Next Generation funds.
@@ B.5 Automatic $\mbox{O}(a)$-improvement and the partially quenched mixed-action setup
We start by considering the argument for automatic $\mbox{O}(a)$-improvement at maximal twist, first put forward in Ref. [57].  [^12] See also Refs. [56, 109, 113, 110, 6] for streamlined versions of the argument and further insights into the structure of automatic $\mbox{O}(a)$-improvement. In the fully twisted case, subtracted standard quark masses are set to zero, possibly up to $\mbox{O}(a)$ effects. Consequently, all terms proportional to $\mathbf{m}$ in Eqs. (B.214)–(B.215) vanish. Equivalently, the terms depending on (combinations of) $m_{i}$ vanish in Eqs. (B.216)–(B.220) and (B.223)–(B.227). Concerning the terms proportional to (combinations of) $\mu_{i}$, it is straightforward to verify that they systematically possess opposite physical parity to the dimension-three fields they correct. While parity is not preserved when $\boldsymbol{\mu}\neq 0$, it is broken only at $\mbox{O}(a)$, which means that the matrix elements of parity-odd operators will be $\mbox{O}(a)$. As a consequence, the insertion of any of these counterterms in on-shell correlation functions will give rise to terms of $\mbox{O}(a^{2})$, since the $\mbox{O}(a)$ amplitude will be multiplied by the explicit factor of $a$ that lowers the field dimension. It follows that no Symanzik counterterms to the currents are needed in order to obtain physical observables that scale towards the continuum as $\mbox{O}(a^{2})$. An equivalent conclusion, as discussed in Ref. [35], can also be obtained using the Wilson Averaging (WA) procedure introduced in Ref. [57], which involves averaging correlation functions with opposite signs of the Wilson parameter $r_{i}$. The discrete symmetries of the Wtm regularisation described in Sect. B.2 allow the WA to be reformulated as a Mass Average (MA), where correlation functions with opposite signs of the subtracted mass $m_{i}$ are considered. The automatic $\mbox{O}(a)$-improvement of physical observables at maximal twist follows from the limiting case in which the MA is performed with $m_{i}=0$ [35]. The argument remains valid when the Sheikholeslami-Wohlert term [108] is included, as its contribution to physical observables in the fully twisted case is of $\mbox{O}(a^{2})$. Moreover, employing a non-perturbative determination of the improvement coefficient $c_{\mathrm{sw}}$ [88, 34] eliminates double insertions of the Pauli term, also contributing at $\mbox{O}(a^{2})$, within the SymEFT framework.
@@
We now turn to the considered mixed-action formulation, corresponding to a partially quenched setup in which the twisted-mass parameters appear exclusively in the valence sector. A first observation can be made when considering quenching. If contributions from sea quarks to the dynamics are neglected, then the terms in Eqs. (B.214)–(B.215) originating from sea-quark loops vanish. These terms can be readily identified as those containing factors of ${\mathrm{tr}}(\mathbf{m})$ and ${\mathrm{tr}}(\boldsymbol{\mu})$, i.e., those associated with improvement coefficients $\bar{b}_{\scriptscriptstyle\Gamma}$ and $\hat{b}_{\scriptscriptstyle\Gamma}$, respectively. For example, in the case of quenched standard Wilson fermions, only counterterms with coefficients $c_{\Gamma}$, $b_{\scriptscriptstyle\Gamma}$, and $f_{\scriptscriptstyle\Gamma}$ need to be considered. Conversely, when present, the terms generated solely by quark loops have coefficients that, in a perturbative expansion, start at $\mbox{O}(\alpha_{\rm\scriptscriptstyle s}^{2}(1/a))$. This applies to both coefficients $\bar{b}_{\scriptscriptstyle\Gamma}$ and $\hat{b}_{\scriptscriptstyle\Gamma}$.
@@
As noted above, the $\mbox{O}(a)$ terms that are not eliminated depend on the improvement coefficients $\bar{b}_{\Gamma}$ which arise only at $\mbox{O}(\alpha_{\rm\scriptscriptstyle s}^{2})$. Moreover, for the considered $N_{\mathrm{\scriptstyle f}}=2+1$ CLS ensembles, the light and strange sea-quark masses satisfy $a\mathrm{tr}\left(\mathbf{m}^{\mathrm{\scriptscriptstyle(sea)}}\right)\lesssim 10^{-2}$. Overall, the $a\bar{b}_{\Gamma}\mathrm{tr}\left(\mathbf{m}^{\mathrm{\scriptscriptstyle(sea)}}\right)$ term is therefore expected to introduce only a very small correction. In the present study, this type of residual contributions are neglected in both the unitary and mixed-action setups. The dedicated numerical studies reported in Sec. 4 indicate that the continuum-limit scaling of the physical quantities based on these two regularisations are consistent with leading lattice artefacts of $\mbox{O}(a^{2})$, as can be seen, for instance, in Fig. 8.
@@ E.1 Statistical error analysis
The choice of $W$ necessitates careful consideration [124, 105]. Values that are too small risk introducing systematic uncertainties through truncation of the tail of the autocorrelation function, whereas excessively large values can lead to inflated statistical fluctuations by incorporating data from regions in which the signal is lost. The exponential decay of $\overline{\Gamma}_{F}^{\alpha}(t)$ at asymptotically large $t$ is governed by the exponential autocorrelation time $\tau_{\mathrm{exp}}^{\alpha}$, corresponding to the slowest mode of the Markov process. The corresponding truncation error is therefore of $\mbox{O}(\exp(-W/\tau_{\mathrm{exp}}^{\alpha}))$. Following the proposal of Ref. [124], we select the value $W=\overline{W}$ that minimises the combined statistical and systematic error, under the assumption $\tau_{\mathrm{exp}}^{\alpha}\approx S_{\tau}\tau_{\mathrm{int}}^{\alpha}$. The default value of the parameter $S_{\tau}$ is set to $S_{\tau}=4$. Direct inspection of the stability of $\tau_{\mathrm{int}}^{\alpha}(F)$ as a function of $W$ is used to adjust the value of $S_{\tau}$ until $\overline{W}$ is observed to lie within a plateau regime.
@@ I Variation in chi-squared weights and model-selection criteria
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1441(8)(4)
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1440(6)(4)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1443(9)(4)
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1441(5)(4)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1441(8)(4)[9]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1440(6)(4)[7]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1441(8)(7)[11]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1440(6)(3)[7]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Wtm}$ 0.1443(9)(4)[10]
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1441(5)(4)[6]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}\;[\mathrm{fm}]\;\mathrm{Combined}$ 0.1440[10]
@@
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1435(8)(4)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (J.288)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1439(8)(4)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (J.289)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1438(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (J.290)
@@
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1445(9)(6)\,\mathrm{fm}\,$, $\mathrm{Wilson}\,$, (J.292)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1449(9)(5)\,\mathrm{fm}\,$, $\mathrm{Wtm}\,$, (J.293)
$\sqrt{t_{0}^{\mathrm{ph}}}$ $=0.1448(6)(4)\,\mathrm{fm}\,$, $\mathrm{Combined}\,.$ (J.294)
@@
Figure 33: Dependence of the value of $\sqrt{t_{0}^{\mathrm{ph}}}$ on the prescription adopted to define the isoQCD limit for the three analysis variants: the unitary Wilson regularisation, the Wtm mixed-action setup, and the combination of both. The prescription adopted in this work follows the Edinburgh consensus, as defined in the FLAG 24 report [9], and presented in Eqs. (4.83)–(4.85), supplemented by the isoQCD value of $f_{K}^{\mathrm{isoQCD}}$ [9], as given in Eq. (4.86). The alternative schemes in Eqs. (J) and (J) are based on the FLAG 21 [8] and FLAG 16 [5] reports, respectively. The $N_{\mathrm{\scriptstyle f}}=2+1$ FLAG 24 average result [9] is denoted by the black filled square symbol and the vertical band.
The results obtained with different isoQCD prescriptions can be mapped onto one another by using the derivatives of the relevant observables with respect to the scheme-defining quantities. Table 10 lists the partial derivatives of $\sqrt{t_{0}^{\mathrm{ph}}}$ with respect to $m_{\pi,K}$ and $f_{\pi,K}$, evaluated at their isoQCD values specified in Eqs. (4.83)–(4.86). We have explicitly verified that the shifts obtained from the derivatives in Table 10 reproduce the values quoted in Eqs. (J.288) and (J.294) – i.e. the results for the FLAG 21 and FLAG 16 isoQCD prescriptions, respectively – when starting from Eq. (4.118) which is based on the FLAG 24 input values in Eqs. (4.83)–(4.86). For related analyses of the scheme dependence in defining the isoQCD limit — for quantities relevant to studies of the hadronic vacuum polarisation contribution to the muon anomalous magnetic moment — we refer to Ref. [4].